OpenStudy (anonymous):

What is 43 base 5 -23 base 5=?

OpenStudy (campbell_st):

43 base 5 is \[4 \times 5^1 + 3 \times 5^0 = 23\] and \[2 \times 5^1 + 3\times 5^0 = 13\] so in a base 10 number system, its 23 - 13 so in base 5 its 43 - 23 ------ 20 base 5 which would be \[2 \times 5^1 + 0 \times 5^0 = 10\] hope it helps

OpenStudy (campbell_st):

a base 5 number system only has 5 digits 0, 1, 2, 3, 4